### v1

@@ Abstract
Let $R$ be a commutative ring with identity and $S\subseteq R$ be a multiplicative set. An ideal $Q$ of $R$ (disjoint from $S$) is said to be $S$-primary if there exists an $s\in S$ such that for all $x,y\in R$ with $xy\in Q$, we have $sx\in Q$ or $sy\in rad(Q)$. Also, we say that an ideal of $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as finite intersection of $S$-primary ideals. In this paper, first we provide an example of $S$-Noetherian ring in which an ideal does not have a primary decomposition. Then our main aim of this paper is to establish the existence and uniqueness of $S$-primary decomposition in $S$-Noetherian rings as an extension of a historical theorem of Lasker-Noether.
  Keywords: $S$-Noetherian ring, $S$-primary ideal, $S$-irreducible ideal, S-primary decompostion.
MSC(2020): 13A15, 13B02, 13C05, 13E05
@@ 1 Introduction
The theory of Noetherian rings played an important role in the development of structure theory of commutative rings. One of the roots of this theory is the historical article [15] by Noether in 1921. Recall that a ring is called Noetherian if it satisfies ascending chain condition on ideals. In the past few decades, several generalizations of Noetherian rings have been extensively studied by many authors because of its importance (see [1], [3], [4], [5], [8] and [12]). As one of its crucial generalizations, Anderson and Dumitrescu [5] introduced $S$-Noetherian rings. A commutative ring $R$ with identity is called $S$-Noetherian, where $S\subseteq R$ is a given multiplicative set, if for each ideal $I$ of $R$, $sI\subseteq J\subseteq I$ for some $s\in S$ and some finitely generated ideal $J$. Theory of primary decomposition, considered as a generalization of the factorization of an integer $n\in\mathbb{Z}$ into a product of prime powers, initiated by Lasker-Noether [11, 15] in their abstract treatment of commutative rings. Lasker-Noether proved that in a commutative Noetherian ring, every ideal can be decomposed as an intersection, called primary decomposition, of finitely many primary ideals (popularly known as the Lasker-Noether decomposition theorem). Due to its segnificance, this theory quickly grew as one of the basic tools of commutative algebra and algebraic geometry. It gives an algebraic foundation for decomposing an algebraic variety into its irreducible components. Recently in [14], Massaoud introduced the concept of $S$-primary ideals as a proper generalization of primary ideals. Let $R$ be a commutative ring with identity and $S\subseteq R$ a multiplicative set. A proper ideal $Q$ (disjoint from $S$) of $R$ is said to be $S$-primary if there exists an $s\in S$ such that for all $a,b\in R$ if $ab\in Q$, then $sa\in Q$ or $sb\in rad(Q)$. The author [14] investigated several properties of this class of ideals and showed that $S$-primary ideals enjoy analogue of many properties of primary ideals. Given the significance of primary decomposition in Noetherian rings, a natural question arises: can we extend the concept of primary decomposition from Noetherian rings to $S$-Noetherian rings ? To extend primary decomposition from Noetherian rings to the broader class of $S$-Noetherian rings, a natural approach is to replace "primary ideals" by "$S$-primary ideals" in the decomposition. By doing this, we can logically adapt this powerful concept that allows us to extend variuos structural properties of Noetherian rings to $S$-Noetherian rings.
@@
In this paper, we introduce the concept of $S$-primary decomposition as a generalization of primary decomposition. We say that an ideal (disjoint with $S$) of a ring $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals of $R$. First we provide an example of $S$-Noetherian ring in which primary decomposition does not exist (see Example 1) which assert that an $S$-Noetherian ring need not be a Laskerian ring in general. Then as one of our main results, we establish the existence of $S$-primary decomposition in $S$-Noetherian rings as a generalization of historical Lasker-Noether decomposition theorem (see Theorem 10 and Theorem 11). Among the other results, we extend first and second uniqueness theorems of primary decomposition to $S$-primary decomposition (see Theorem 16 and Theorem 19).
@@ 2 $S$-Primary Decomposition in $S$-Noetherian Ring
It is well known that primary decomposition exists in a Noetherian ring. Recall from [1] that a ring $R$ is said to have a Noetherian spectrum if $R$ satisfies the ascending chain condition (ACC) on radical ideals. This is equivalent to the condition that $R$ satisfies the ACC on prime ideals, and each ideal has only finitely many prime ideals minimal over it. Also, a ring $R$ is said to be Laskerian if each ideal of $R$ has a primary decomposition. We begin by providing an example of an $S$-Noetherian ring in which primary decomposition does not hold.
@@
Let $R=F[x_{1},x_{2},\ldots,x_{n},\ldots]$ be a polynomial ring in infinitely many indeterminates over a field $F$. Since $R$ has an ascending chain of prime ideals $(x_{1})\subseteq(x_{1},x_{2})\subseteq\cdots\subseteq(x_{1},x_{2},\ldots,x_{n})\subseteq\cdots$ which does not terminate, so $R$ has no Noetherian spectrum. This implies that $R$ is a non-Laskerian ring since every Laskerian ring has a Noetherian spectrum (see [10, Theorem 4]). Consider a multiplicative set $S=R\setminus\{0\}$. Then by [5, Proposition 2(a)], $R$ is an $S$-Noetherian ring. Hence $R$ is an $S$-Noetherian ring but not Laskerian.
Thus we need $S$-version of primary decomposition of ideals. Now we define the concept of $S$-primary decomposition of ideals as a generalization of primary decomposition.
@@
An ideal $Q$ (disjoint with $S$) of a ring $R$ is said to be $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals of $R$. Further, an $S$-primary decomposition $I=\bigcap\limits_{i=1}^{n}Q_{i}$ is said to be minimal if the $rad(Q_{i})$ are all distinct, and $\bigcap\limits_{j\neq i}Q_{j}\nsubseteq Q_{i}$ for all $i=1,2,\ldots,n$.
It is clear from the definition that the concepts of $S$-primary decomposition and primary decomposition coincide for $S=\{1\}$. The following examples show that the concept of $S$-primary decomposition is a proper generalization of the concept of primary decomposition.
@@
[…]
Example 4.
Consider the Boolean ring $R=\mathcal{P}(\mathbb{N})$, the power set of $\mathbb{N}$, where addition in $R$ being the symmetric difference of sets, and multiplication being the intersection of sets. By [17, Theorem 1], the zero ideal $(0)$ in $R$ does not have a primary decomposition. Clearly, $R$ is not Noetherian. Consider a multiplicative set $S=R\setminus\{0\}$. Then by [5, Proposition 2(a)], $R$ is an $S$-Noetherian ring. Next, we show that $(0)$ is an $S$-primary ideal of $R$. For this, let $A,B\in R$ such that $AB=0$ and $sA\neq 0$ for all $s\in S$. This implies that $A$ is a non-zero element of $R$, and hence $A\in S$. Put $s=A$. Then we have $sB=0$, as desired. Hence $(0)$ is $S$-primary decomposable.
Recall from [7], an ideal $I$ of the ring $R$ is called irreducible if $I=J\cap K$ for some ideals $J$, $K$ of $R$, then either $I=J$ or $I=K$. It is well known that the classical proof of existence of primary decomposition in Noetherian ring involves the concept of irreducible ideals. So we need $S$-version of irreducible ideals to prove the existence of $S$-primary decomposition in $S$-Noetherian rings.
@@
An ideal $Q$ (disjoint with $S$) of the ring $R$ is called $S$-irreducible if $s(I\cap J)\subseteq Q\subseteq I\cap J$ for some $s\in S$ and some ideals $I$, $J$ of $R$, then there exists $s^{\prime}\in S$ such that either $ss^{\prime}I\subseteq Q$ or $ss^{\prime}J\subseteq Q$.
@@
Let $R=F[x,y,z]$, where $F$ is a field, and $S=\{z^{k}\hskip 2.84544pt|\hskip 2.84544ptk\in\mathbb{N}\cup\{0\}\}$ be a multiplicative set. Since $R$ is a Noetherian ring, so $R$ is an $S$-Noetherian ring. Then the ideal $Q=(x,yz)$ is not irreducible since $Q=(x,y)\cap(x,z)$. On the other hand if we take $s=z\in S$, then $s(x,y)=(sx,sy)\subseteq Q$. Thus $Q$ is an $S$-irreducible ideal of $R$.
Recall from [2], an ideal $Q$ (disjoint with $S$) of a ring $R$ is said to be an $S$-prime if there exists an $s\in S$ such that for $a,b\in R$ with $ab\in Q$, we have either $sa\in Q$ or $sb\in Q$. Clearly every $S$-prime ideal is $S$-primary. The following example shows that converse of this is not true in general.
@@
Consider $R=\mathbb{Z}$, $Q=4\mathbb{Z}$, and $S=\mathbb{Z}\setminus 2\mathbb{Z}$. Notice that $4\in Q$ but $2s\notin Q$ for all $s\in S$. This implies that $Q$ is not an $S$-prime ideal of $R$. Now we prove that $Q$ is an $S$-primary ideal of $R$. For this, let $a,b\in R$ such that $ab\in Q$. Then we have two cases:
Case-1 If $a=2r$ and $b=2k$, where $r$, $k\in\mathbb{Z}$, then $ab\in Q$. In this case, there always exists $s\in S$ such that either $sa\in rad(Q)$ or $sb\in rad(Q)$. Thus $Q$ is $S$-primary.
Case-2 If either $a=4m$ or $b=4n$, where $m$, $n\in\mathbb{Z}$, then $ab\in Q$. Also in this case, there always exists $s\in S$ such that either $sa\in rad(Q)$ or $sb\in rad(Q)$. Therefore $Q$ is a $S$-primary ideal of $R$.
Recall from [14] that if $Q$ is an $S$-primary ideal of a ring $R$ and $rad(Q)=P$ is an $S$-prime ideal, then $Q$ is said to be an $S$-$P$-primary ideal of $R$.
@@
Finite intersection of $S$-$P$-primary ideals is $S$-$P$-primary.
@@
Let $Q_{1},Q_{2},\ldots,Q_{n}$ be $S$-$P$-primary ideals, then $S\cap Q_{i}=\emptyset$ for each $i=1,2,\ldots,n$, and so $S\cap(\bigcap\limits_{i=1}^{n}Q_{i})=\emptyset$. Suppose $Q=\bigcap\limits_{i=1}^{n}Q_{i}$. Since each $Q_{i}$ is $S$-$P$-primary, $rad(Q)=rad(\bigcap\limits_{i=1}^{n}Q_{i})=\bigcap\limits_{i=1}^{n}rad(Q_{i})=P$. Now let $xy\in Q$ with $sy\notin Q$ for all $s\in S$. Consequently, there exists $k$, $(1\leq k\leq n)$ such that $xy\in Q_{k}$ and $sy\notin Q_{k}$ for all $s\in S$. This implies that there exists $s\in S$ such that $sx\in rad({Q_{k}})=P$. Hence $Q$ is $S$-$P$-primary. ∎
In view of the Proposition 8, we can always obtain minimal $S$-primary decomosition from $S$-primary decomosition by replacing all $S$-primary ideals with the same radical by their intersection. Following [6], let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $\mathcal{I}$ (disjoint with $S$) be an ideal of $R$. Then $\mathcal{I}$ is said to be $S$-maximal if there exists $s\in S$ such that whenever $\mathcal{I}\subseteq J$ for some ideal $J$ of $R$, then either $sJ\subseteq\mathcal{I}$ or $J\cap S\neq\emptyset$.
Proposition 9.
Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $Q$ (disjoint with $S$) be an ideal of $R$. If $rad(Q)=\mathcal{M}$ is $S$-maximal, then $Q$ is an $S$-primary ideal of $R$.
Proof.
Let $a,b\in R$ such that $ab\in Q$ and $sb\notin rad{(Q)}=\mathcal{M}$ for all $s\in S$. Choose a maximal ideal $\mathcal{M^{\prime}}$ of $R$ such that $\mathcal{M}\subseteq\mathcal{M^{\prime}}$ and $\mathcal{M^{\prime}}\cap S=\emptyset$. This implies that there exists $s^{\prime}\in S$ such that $s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}$ since $\mathcal{M}$ is $S$-maximal. Consequently, $sb\notin s^{\prime}\mathcal{M^{\prime}}$ since $sb\notin\mathcal{M}$ for all $s\in S$. If $b\in\mathcal{M^{\prime}}$, then $s^{\prime}b\in s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}$, a contradiction since $sb\notin\mathcal{M}$ for all $s\in S$. Therefore $b\notin\mathcal{M^{\prime}}$. This implies that $\mathcal{M^{\prime}}+Rb=R$. Write $1=c+\lambda b$ for some $c\in\mathcal{M^{\prime}}$ and $\lambda\in R$, then $s^{\prime}=s^{\prime}c+\lambda s^{\prime}b$. Now since $s^{\prime}c\in s^{\prime}\mathcal{M^{\prime}}\subseteq\mathcal{M}=rad(Q)$, there exists $k\geq 1$ such that $(s^{\prime}c)^{k}\in Q$. Then by binomial theorem $(s^{\prime})^{k}=(s^{\prime}c+\lambda s^{\prime}b)^{k}=\binom{k}{0}(s^{\prime}c)^{k}+\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda s^{\prime}b)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda s^{\prime}b)^{2}+\cdots+\binom{k}{k}(s^{\prime}c)^{k-k}(\lambda s^{\prime}b)^{k}=(s^{\prime}c)^{k}+s^{\prime}b\left(\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda)^{2}(s^{\prime}b)+\cdots+(\lambda)^{k}(s^{\prime}b)^{k-1}\right)$=$(s^{\prime}c)^{k}+\lambda^{\prime}s^{\prime}b$ where $\lambda^{\prime}=\binom{k}{1}(s^{\prime}c)^{k-1}(\lambda)+\binom{k}{2}(s^{\prime}c)^{k-2}(\lambda)^{2}(s^{\prime}b)+\cdots+(\lambda)^{k}(s^{\prime}b)^{k-1}\in R$. Then $(s^{\prime})^{k}a=a(s^{\prime}c)^{k}+\lambda^{\prime}s^{\prime}ab$, and so $(s^{\prime})^{k}a\in Q$ since $(s^{\prime}c)^{k}\in Q$ and $ab\in Q$. Consequently, $(s^{\prime}a)^{k}=(s^{\prime})^{k}a(a)^{k-1}\in Q$ since $(s^{\prime})^{k}a\in Q$. Hence $Q$ is an $S$-primary ideal of $R$. ∎
Following [3], let $E$ be a family of ideals of a ring $R$. An element $I\in E$ is said to be an $S$-maximal element of $E$ if there exists an $s\in S$ such that for each $J\in E$, if $I\subseteq J$, then $sJ\subseteq I$. Also an increasing sequence $(I_{i})_{i\in\mathbb{N}}$ of ideals of $R$ is called $S$-stationary if there exists a positive integer $k$ and $s\in S$ such that $sI_{n}\subseteq I_{k}$ for all $n\geq k$.
@@
Suppose $Q$ is an $S$-irreducible ideal of $R$. Let $a,b\in R$ be such that $ab\in Q$ and $sb\notin Q$ for all $s\in S$. Our aim is to show that there exists $t\in S$ such that $ta\in rad(Q).$ Consider $A_{n}=\{x\in R\hskip 5.69046pt|\hskip 2.84544pta^{n}x\in Q\}$ for $n\in\mathbb{N}.$ Clearly each $A_{n}$ is an ideal of $R$ and $A_{1}\subseteq A_{2}\subseteq A_{3}\subseteq\cdots$ is an increasing chain of ideals of $R$. Since $R$ is $S$-Noetherian, then by [9, Theorem 2.3], this chain is $S$-stationary, i.e., there exist $k\in\mathbb{N}$ and $s\in S$ such that $sA_{n}\subseteq A_{k}$ for all $n\geq k$. Consider the two ideals $I=\textless a^{k}\textgreater+\hskip 2.84544ptQ$ and $J=\textless b\textgreater+\hskip 2.84544ptQ$ of $R$. Then $Q\subseteq I\cap J$. For the reverse containment, let $y\in I\cap J$. Write $y=a^{k}z+q$ for some $z\in R$ and $q\in Q$. Since $ab\in Q$, $aJ\subseteq Q$; whence $ay\in Q$. Now $a^{k+1}z=a(a^{k}z)=a(y-q)\in Q$. This implies that $z\in A_{k+1}$, and so $sz\in sA_{k+1}\subseteq A_{k}$. Consequently, $a^{k}sz\in Q$ which implies that $a^{k}sz+sq=sy\in Q$. Thus we have $s(I\cap J)\subseteq Q\subseteq I\cap J$. This implies that there exists $s^{\prime}\in S$ such that either $ss^{\prime}I\subseteq Q$ or $ss^{\prime}J\subseteq Q$ since $Q$ is $S$-irreducible. If $ss^{\prime}J\subseteq Q$, then $ss^{\prime}b\in Q$ which is not possible. Therefore $ss^{\prime}I\subseteq Q$ which implies that $ss^{\prime}a^{k}\in Q$. Put $t=ss^{\prime}\in S$. Then $(ta)^{k}\in Q$ and hence $ta\in rad(Q)$, as desired. ∎

### v2

@@ Abstract
Let $R$ be a commutative ring with identity, and let $S\subseteq R$ be a multiplicative set. An ideal $Q$ of $R$ (disjoint from $S$) is said to be $S$-primary if there exists an $s\in S$ such that for all $x,y\in R$ with $xy\in Q$, we have $sx\in Q$ or $sy\in rad(Q)$. Also, we say that an ideal of $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals. In this paper, first we provide an example of an $S$-Noetherian ring in which an ideal does not have a primary decomposition. Then our main aim of this paper is to establish the existence and uniqueness of $S$-primary decomposition in $S$-Noetherian rings as an extension of a historical theorem of Lasker-Noether.
Keywords: $S$-Noetherian ring, $S$-primary ideal, $S$-irreducible ideal, S-primary decomposition.
MSC(2010): 13C05, 13B02, 13E05.
@@ 1 Introduction
The theory of Noetherian rings has been playing an important role in the development of structure theory of commutative rings. One of the roots of this theory is the historical article [14] by Noether in 1921. Recall that a ring is called Noetherian if it satisfies ascending chain condition on ideals. In the past few decades, several generalizations of Noetherian rings have been extensively studied by many authors because of its importance (see [1], [3], [4], [5], [7], and [12]). As one of its crucial generalizations, Anderson and Dumitrescu [5] introduced $S$-Noetherian rings. A commutative ring $R$ with identity is called $S$-Noetherian, where $S\subseteq R$ is a given multiplicative set, if for each ideal $I$ of $R$, $sI\subseteq J\subseteq I$ for some $s\in S$ and some finitely generated ideal $J$. Theory of primary decomposition, considered as a generalization of the factorization of an integer $n\in\mathbb{Z}$ into a product of prime powers, initiated by Lasker-Noether [11, 14] in their abstract treatment of commutative rings. Lasker-Noether proved that in a commutative Noetherian ring, every ideal can be decomposed as an intersection, called primary decomposition, of finitely many primary ideals (popularly known as the Lasker-Noether decomposition theorem). Due to its significance, this theory quickly grew as one of the basic tools of commutative algebra and algebraic geometry. It gives an algebraic foundation for decomposing an algebraic variety into its irreducible components. Recently in [13], Massaoud introduced the concept of $S$-primary ideals as a proper generalization of primary ideals. Let $R$ be a commutative ring with identity and $S\subseteq R$ a multiplicative set. A proper ideal $Q$ (disjoint from $S$) of $R$ is said to be $S$-primary if there exists an $s\in S$ such that for all $a,b\in R$ if $ab\in Q$, then $sa\in Q$ or $sb\in rad(Q)$. The author [13] investigated several properties of this class of ideals and showed that $S$-primary ideals enjoy analogue of many properties of primary ideals. Given the significance of primary decomposition in Noetherian rings, a natural question arises:
Question 1.1.
Can the idea of primary decomposition in Noetherian rings be extended to $S$-Noetherian rings?
We provide a positive answer to the above question in this paper. A natural way to extend primary decomposition from Noetherian rings to the broader class of $S$-Noetherian rings is to replace "primary ideals" by "$S$-primary ideals" in the decomposition process. By doing this, we can logically adapt this powerful concept that allows us to extend various structural properties of Noetherian rings to $S$-Noetherian rings.
@@
In this paper, we introduce the concept of $S$-primary decomposition as a generalization of primary decomposition. We say that an ideal (disjoint from $S$) of a ring $R$ is $S$-primary decomposable or has an $S$-primary decomposition if it can be written as a finite intersection of $S$-primary ideals of $R$. First we provide an example of an $S$-Noetherian ring in which primary decomposition does not exist (see Example 1) which asserts that an $S$-Noetherian ring need not be a Laskerian ring in general. Then as one of our main results, we establish the existence of $S$-primary decomposition in $S$-Noetherian rings as a generalization of historical Lasker-Noether decomposition theorem (see Theorem 8 and Theorem 9). Among the other results, we extend first and second uniqueness theorems of primary decomposition to $S$-primary decomposition (see Theorem 16 and Theorem 19).
@@ 2 $S$-Primary Decomposition in $S$-Noetherian Ring
It is well known that primary decomposition exists in a Noetherian ring. Recall from [1] that a ring $R$ is said to have a Noetherian spectrum if $R$ satisfies the ascending chain condition (ACC) on radical ideals. This is equivalent to the condition that $R$ satisfies the ACC on prime ideals, and each ideal has only finitely many prime ideals minimal over it. Also, a ring $R$ is said to be Laskerian if each ideal of $R$ has a primary decomposition.
Recall that [5], an ideal $I$ of $R$ is called $S$-finte if $sI\subseteq J\subseteq I$ for some finitely generated ideal $J$ of $R$ and some $s\in S$. Then $R$ is said to be an $S$-Noetherian ring if each ideal of $R$ is $S$-finite. We begin by providing an example of an $S$-Noetherian ring in which primary decomposition does not hold.
@@
Let $R=F[x_{1},x_{2},\ldots,x_{n},\ldots]$ be the polynomial ring in infinitely many indeterminates over a field $F$. Since $R$ has an ascending chain of prime ideals $(x_{1})\subseteq(x_{1},x_{2})\subseteq\cdots\subseteq(x_{1},x_{2},\ldots,x_{n})\subseteq\cdots$ which does not terminate, so $R$ has no Noetherian spectrum. This implies that $R$ is a non-Laskerian ring, by [9, Theorem 4]. Consider the multiplicative set $S=R\setminus\{0\}$. Then by [5, Proposition 2(a)], $R$ is an $S$-Noetherian ring. Hence $R$ is an $S$-Noetherian ring but not Laskerian.
Recall that let $f:R\longrightarrow S^{-1}R$ denote the usual homomorphism of rings given by $f(r)=\frac{r}{1}$. For any ideal $I$ of $R$, the $f^{-1}(S^{-1}I)$ called the contraction of $I$ with respect to $S$, that is, $\{a\in R\hskip 5.69046pt|\hskip 5.69046pt\dfrac{a}{1}\in S^{-1}I\}$ is denoted by $S(I)$. Notice that $I\subseteq S(I)$. Thus we need an $S$-version of primary decomposition of ideals. Now we define the concept of $S$-primary decomposition of ideals as a generalization of primary decomposition.
@@
Let $R$ be a ring and let $S$ be a multiplicative set of $R$. Let $I$ be an ideal of $R$ such that $I\cap S=\emptyset$. We say that $I$ admits an $S$-primary decomposition if $I$ is a finite intersection of $S$-primary ideals of $R$. In such a case, we say that $I$ is $S$-decomposable. An $S$-primary decomposition $I=\bigcap\limits_{i=1}^{n}Q_{i}$ of $I$ with $rad(Q_{i})=P_{i}$ for each $i\in\{1,2,\ldots,n\}$ is said to be minimal if the following conditions hold:
  1. 1.
    $S(P_{i})\neq S(P_{j})$ for all distinct $i,j\in\{1,2,\ldots,n\}$.
  2. 2.
    $S(Q_{i})\nsupseteq\bigcap_{j\in\{1,2,\ldots,n\}\setminus\{i\}}S(Q_{j})$ for each $i\in\{1,2,\ldots,n\}$ (equivalently,
    $S(Q_{i})\nsupseteq\bigcap_{j\in\{1,2,\ldots,n\}\setminus\{i\}}Q_{j}$ for each $i\in\{1,2,\ldots,n\})$.
From definition 2, the concepts of $S$-primary decomposition and primary decomposition coincide for $S=\{1\}$. The following example shows that the concept of $S$-primary decomposition is a proper generalization of the concept of primary decomposition.
@@
Consider the Boolean ring $R=\prod_{n=1}^{\infty}\mathbb{Z}_{2}$ (countably infinite copies of $\mathbb{Z}_{2})$. According to [16, Theorem 1], the zero ideal $(0)=(0,0,0,\ldots)$ in $R$ does not have the primary decomposition. Clearly, $R$ is not Noetherian. Consider the multiplicative set $S=\{{1_{R}}=(1,1,1,\ldots),s=(1,0,0,\ldots)\}$. Let $I$ be an ideal of $R$. Either $s\in I$ or $s\notin I$. If $s\in I$, then $sI\subseteq Rs\subseteq I$ and as $Rs$ is a finitely generated ideal of $R$, it follows that $I$ is $S$-finite. If $s\notin I$, then $I\subseteq(0)\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\cdots$. In such a case, $sI\subseteq(0)\times(0)\times(0)\times\cdots\subseteq I$. Hence, $I$ is $S$-finite. This shows that $R$ is an $S$-Noetherian ring. Next, we show that $(0)$ is an $S$-primary ideal of $R$. First, we observe that $(0)\cap S=\emptyset$. Now, let $a=(a_{n})_{n\in\mathbb{N}}$, $b=(b_{n})_{n\in\mathbb{N}}\in R$ such that $ab=0$, where each $a_{i},{b_{i}}\in\mathbb{Z}_{2}$. This implies that $a_{n}b_{n}=0$ for all $n\in\mathbb{N}$, in particular, $a_{1}b_{1}=0$. Then we have either $a_{1}=0$ or $b_{1}=0$. If $a_{1}=0$, then $sa=0$. If $b_{1}=0$, then $sb=0$. Thus $(0)$ is an $S$-primary ideal. Therefore $(0)$ is $S$-primary decomposable.
Recall from [6], an ideal $I$ of the ring $R$ is called irreducible if $I=J\cap K$ for some ideals $J$, $K$ of $R$, then either $I=J$ or $I=K$. It is well known that the classical proof of existence of primary decomposition in a Noetherian ring involves the concept of irreducible ideals. So we need $S$-version of irreducible ideals to prove the existence of $S$-primary decomposition in $S$-Noetherian rings.
@@
An ideal $Q$ (disjoint from $S$) of the ring $R$ is called $S$-irreducible if $s(I\cap J)\subseteq Q\subseteq I\cap J$ for some $s\in S$ and some ideals $I$, $J$ of $R$, then there exists $s^{\prime}\in S$ such that either $ss^{\prime}I\subseteq Q$ or $ss^{\prime}J\subseteq Q$.
@@
Let $R=\mathbb{Z}$, $S=\mathbb{Z}\setminus 3\mathbb{Z}$ and $I=6\mathbb{Z}$. Since $I=2\mathbb{Z}\cap 3\mathbb{Z}$, therefore $I$ is not an irreducible ideal of $R$. Now take $s=2\in S$. Then $2(3\mathbb{Z})=6\mathbb{Z}\subseteq I$. Thus $I$ is an $S$-irreducible ideal of $R$.
Recall from [2], an ideal $Q$ (disjoint from $S$) of a ring $R$ is said to be $S$-prime if there exists an $s\in S$ such that for $a,b\in R$ with $ab\in Q$, we have either $sa\in Q$ or $sb\in Q$. Clearly, every $S$-prime ideal is $S$-primary. The following example shows that the converse of this is not true in general.
@@
Consider $R=\mathbb{Z}$, $Q=4\mathbb{Z}$, and $S=\mathbb{Z}\setminus 2\mathbb{Z}$. Notice that $4\in Q$ but $2s\notin Q$ for all $s\in S$. This implies that $Q$ is not an $S$-prime ideal of $R$. Obviously, $Q$ is a primary ideal of $R$ and hence $S$-primary.
Recall from [13, Proposition 2.5] that if $Q$ is an $S$-primary ideal of a ring $R$, then $P=\text{rad}(Q)$ is an $S$-prime ideal. In such a case, we say that $Q$ is an $P$-$S$-primary ideal of $R$.
@@
Let $S$ be a multiplicative set of a ring $R$. Then the following statements hold:
  1. 1.
    Finite intersection of $P$-$S$-primary ideals is $P$-$S$-primary.
  2. 2.
    If $Q$ is a $P$-primary ideal of $R$ with $Q\cap S=\emptyset$, then for any ideal $J$ of $R$ with $J\cap S\neq\emptyset$, $Q\cap J$ is a $(P\cap rad(J))$-$S$-primary ideal of $R$.
@@
  1. 1.
    Let $Q_{1},Q_{2},\ldots,Q_{n}$ be $P$-$S$-primary ideals, then $S\cap Q_{i}=\emptyset$ for each $i=1,2,\ldots,n$, and so $S\cap(\bigcap\limits_{i=1}^{n}Q_{i})=\emptyset$. Suppose $Q=\bigcap\limits_{i=1}^{n}Q_{i}$. Since each $Q_{i}$ is $P$-$S$-primary, $rad(Q)=rad(\bigcap\limits_{i=1}^{n}Q_{i})=\bigcap\limits_{i=1}^{n}rad(Q_{i})=P$. Now, let $xy\in Q$, where $x,y\in R$ and with $sy\notin Q$ for all $s\in S$. Consequently, for every $s\in S$, there exists $k_{s}$ such that $xy\in Q_{k_{s}}$ and $sy\notin Q_{k_{s}}$. Let $s_{i}\in S$ be the element satisfying the $S$-primary property for $Q_{i}$. Since we have finitely many $Q_{i}$, put $s=s_{1}s_{2}\dots s_{n}\in S$. Now fix $s$ and assume that $xy\in Q$ but $sy\notin Q$. Thus, there exists $k$ such that $xy\in Q_{k}$ and $sy\notin Q_{k}$. Then for $s_{k}\in S$ we obtain $s_{k}x\in rad(Q_{k})=P$ or $s_{k}y\in Q_{k}$. The latter case gives $sy\in Q_{k}$, a contradiction. Thus $sx\in rad(Q)=P$, and therefore $Q$ is $P$-$S$-primary.
  2. 2.
    As $Q\cap S=\emptyset$, it follows that $(Q\cap J)\cap S=\emptyset$. By assumption, $J\cap S\neq\emptyset$. Let $s\in J\cap S$. Let $a,b\in R$ be such that $ab\in Q\cap J$. Either $a\in Q$ or $b\in rad(Q)=P$, since $Q$ is a $P$-primary ideal of $R$. Hence, either $sa\in Q\cap J$ or $sb\in P\cap J\subseteq P\cap rad(J)=rad(Q)\cap rad(J)=rad(Q\cap J)$. This proves that $Q\cap J$ is a $(P\cap rad(J))$-$S$-primary ideal of $R$.
If an ideal $I$ of a ring $R$ admits an $S$-primary decomposition, then we prove in Remark 11 that it will admit a minimal $S$-primary decomposition.
Following [3], let $E$ be a family of ideals of a ring $R$. An element $I\in E$ is said to be an $S$-maximal element of $E$ if there exists an $s\in S$ such that for each $J\in E$, if $I\subseteq J$, then $sJ\subseteq I$. Also an increasing sequence $(I_{i})_{i\in\mathbb{N}}$ of ideals of $R$ is called $S$-stationary if there exist a positive integer $k$ and $s\in S$ such that $sI_{n}\subseteq I_{k}$ for all $n\geq k$.
@@
Suppose $Q$ is an $S$-irreducible ideal of $R$. Let $a,b\in R$ such that $ab\in Q$ and $sb\notin Q$ for all $s\in S$. Our aim is to show that there exists $t\in S$ such that $ta\in rad(Q)$. Consider $A_{n}=\{x\in R\hskip 5.69046pt|\hskip 2.84544pta^{n}x\in Q\}$ for $n\in\mathbb{N}.$ Clearly each $A_{n}$ is an ideal of $R$ and $A_{1}\subseteq A_{2}\subseteq A_{3}\subseteq\cdots$ is an increasing chain of ideals of $R$. Since $R$ is $S$-Noetherian, by [8, Theorem 2.3], this chain is $S$-stationary, i.e., there exist $k\in\mathbb{N}$ and $s\in S$ such that $sA_{n}\subseteq A_{k}$ for all $n\geq k$. Consider the two ideals $I=\textless a^{k}\textgreater+\hskip 2.84544ptQ$ and $J=\textless b\textgreater+\hskip 2.84544ptQ$ of $R$. Then $Q\subseteq I\cap J$. For the reverse containment, let $y\in I\cap J$. Write $y=a^{k}z+q$ for some $z\in R$ and $q\in Q$. Since $ab\in Q$, $aJ\subseteq Q$; whence $ay\in Q$. Now $a^{k+1}z=a(a^{k}z)=a(y-q)\in Q$. This implies that $z\in A_{k+1}$, and so $sz\in sA_{k+1}\subseteq A_{k}$. Consequently, $a^{k}sz\in Q$ which implies that $a^{k}sz+sq=sy\in Q$. Thus we have $s(I\cap J)\subseteq Q\subseteq I\cap J$. This implies that there exists $s^{\prime}\in S$ such that either $ss^{\prime}I\subseteq Q$ or $ss^{\prime}J\subseteq Q$ since $Q$ is $S$-irreducible. If $ss^{\prime}J\subseteq Q$, then $ss^{\prime}b\in Q$ which is not possible. Therefore $ss^{\prime}I\subseteq Q$ which implies that $ss^{\prime}a^{k}\in Q$. Put $t=ss^{\prime}\in S$. Then $(ta)^{k}\in Q$ and hence $ta\in rad(Q)$, as desired. ∎